Circular No. 7406

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000bs IN UGC 10710 M. Papenkova and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery of an apparent supernova (mag ∼ 17.7) on unfiltered images taken on Apr. 19.5 and 20.3 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2000bs is located at α = 17^{h}06^{m}54^{s}.67, δ = +43^{o}06′41′′.7 (equinox 2000.0), which is 23′′.9 east and 38′′.0 south of the nucleus of UGC 10710. An unfiltered image taken on Apr. 8.5 showed nothing at the position of the new object (limiting mag ∼ 19.0).

SUPERNOVAE 2000bt, 2000bu, 2000bv, 2000bw, 2000bx, 2000by, 2000bz

D. Hardin, LPNHE, Paris, reports for N. Walton, on behalf of the EROS collaboration and the European Supernova Cosmology Consortium (cf. IAUC 7258, plus A. Amadon, Service de Physique des Particules, Saclay; D. Batcheldor, University of Hertfordshire; G. Folatelli, G. Garavini, and S. Nobili, University of Stockholm; and A. Mourão and M. J. Cruz, Instituto Superior Técnico, Lisbon), the discovery of seven supernovae in the course of the automated supernova search with the 1-m Marly telescope at the European Southern Observatory (ESO). All were found by subtraction of CCD images from reference images taken in early March (limiting mag ∼ 21.5), and each was within 2′′ of its host galaxy’s center; most were discovered near maximum light, but SN 2000bt was discovered ∼ 3 weeks past maximum. The supernovae were confirmed via images taken 1 or 2 days after discovery at the Marly telescope, except SN 2000bz (which was confirmed at the ESO 3.6-m telescope). Confirming spectroscopy was obtained at the 4.2-m William Herschel Telescope during Apr. 7–9 and at the ESO 3.6-m telescope during Apr. 9–11. Finding charts can be found at

http://supernovae.in2p3.fr/Intz/INT200004/sn.html.

SN  2000 UT  α_{2000}  δ_{2000}  V  z Type
2000bt Mar. 26 10^{h}16^{m}18.^{s}06 − 5^{o}44′47′′.3 19.6 0.04 Ia
2000bu Mar. 31 11 27 11.45 − 6 23 14.6 19.4 0.05 II?
2000bv Apr. 1 12 59 28.70 −12 20 07.6 20.6 0.12 II?
2000bw Apr. 4 11 09 49.85 − 4 24 46.4 20.5 0.12 II?
2000bx Apr. 6 13 48 55.53 − 6 18 35.9 19.2 0.09 Ia
2000by Apr. 7 11 39 54.91 − 4 22 16.4 19.3 0.10 Ia
2000bz Apr. 8 14 15 02.66 − 6 17 16.0 21.2 0.26 Ia

2000 April 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green